# Can LLMs Beat Classical Hyperparameter Optimization Algorithms? A Study on autoresearch

### 25 Mar 2026 v1

@@ Abstract
The autoresearch repository enables an LLM agent to search for optimal hyperparameter configurations on an unconstrained search space by editing the training code directly. Given a fixed compute budget and constraints, we use autoresearch as a testbed to compare classical hyperparameter optimization (HPO) algorithms against LLM-based methods on tuning the hyperparameters of a small language model. Within a fixed hyperparameter search space, classical HPO methods such as CMA-ES and TPE consistently outperform LLM-based agents. However, an LLM agent that directly edits training source code in an unconstrained search space narrows the gap to classical methods substantially despite using only a self-hosted open-weight 27B model. Methods that avoid out-of-memory failures outperform those with higher search diversity, suggesting that reliability matters more than exploration breadth. While small and mid-sized LLMs struggle to track optimization state across trials, classical methods lack domain knowledge. To bridge this gap, we introduce Centaur, a hybrid that shares CMA-ES’s internal state, including mean vector, step-size, and covariance matrix, with an LLM [^1] Named after the mythological half-human, half-horse hybrid: our method merges LLM reasoning with classical optimization.. Centaur achieves the best result in our experiments, with its 0.8B variant outperforming the 27B variant, suggesting that a cheap LLM suffices when paired with a strong classical optimizer. The 0.8B model is insufficient for unconstrained code editing but sufficient for hybrid optimization, while scaling to 27B provides no advantage for fixed search space methods with the open-weight models tested. Code is available at https://github.com/ferreirafabio/autoresearch-automl.
@@ 1 Introduction
To answer these questions, we benchmark 9 HPO methods, spanning 4 classical, 4 LLM-based, and 1 hybrid, on Karpathy’s autoresearch task. All methods operate under the same 24-hour GPU training budget with 3 seeds. To reduce human priors, we automatically extract 14 hyperparameters from the training script; while the ranges require some domain knowledge, the HP selection itself is automated, removing manual search space curation. Figure˜1 compares all 9 methods using 27B LLM variants against cumulative training wall-time, showing that classical methods find better configurations than LLM-based agents within the fixed search space. The exception is Karpathy Agent (Code), which directly edits training source code and is competitive with classical methods, even though the latter find a similar performing hyperparameter configuration ${\sim}4{\times}$ faster. Since all LLM methods use a self-hosted open-weight model (Qwen3.5-27B), stronger frontier models may narrow this gap further.
@@
  • We show that classical HPO outperforms LLM agents within a fixed search space, while an LLM agent editing code directly narrows the gap substantially despite using only a self-hosted open-weight model.
@@
  • We analyze search diversity, OOM rates, and model scaling across all methods, and release per-trial LLM conversation logs.
@@ 3 Experimental Setup
We ran each method for 24 hours with three seeds and report results both against 24-hour wall-clock time and on trial scale (up to 300 trials). Methods with high OOM rates accumulated more trials due to fast failures (<30{<}30s per OOM) but show negligible improvement beyond trial 300. We report failed trials (OOM) as val_bpb=100.0{}=100.0, a finite penalty orders of magnitude worse than any valid result while remaining compatible with all surrogate models, so that optimizers learn to avoid infeasible regions.
All LLM-based methods use Qwen3.5 (Team, 2026) as the LLM optimizer, self-hosted via vLLM (Kwon et al., 2023) on the same GPU that trains the optimizee (the $\sim$}50M-parameter language model). To ensure equal resource allocation, we capped training VRAM to 80 GB for all methods, comparable to the H100 used by Karpathy (Karpathy, 2025a), and reserved the remaining memory for the vLLM server. We disabled Qwen3.5’s thinking mode and sampled with temperature 0.7, limiting outputs to 2048 tokens for fixed-HP methods and 16384 tokens for code editing. All prompt templates are available in our repository.
@@ 4 Centaur: CMA-ES Guided LLM Optimization
Centaur shares CMA-ES’s full internal state with the LLM on a fraction of trials. On every trial, CMA-ES proposes a candidate configuration from its multivariate Gaussian, parameterized by mean vector $\bm{\mu}$, step-size $\sigma$, and covariance matrix $\mathbf{C}$. On 30% of trials, the LLM receives CMA-ES’s proposal along with $\bm{\mu}$, $\sigma$, $\mathbf{C}$, the top-5 configurations, and the last 20 trials. The LLM may override the proposal, and in practice does so in nearly all cases: 100% of the time with the 27B model and 95% with the 0.8B model. Crucially, CMA-ES updates its internal state from all trial results, including those where the LLM overrode its proposal, so the optimizer continuously learns from the full trajectory.
We chose CMA-ES because its internal state is particularly interpretable for LLM communication: the mean is a concrete configuration, $\sigma$ is a single scalar, and $\mathbf{C}$ is a labeled matrix. In contrast, TPE maintains density estimators that are difficult to summarize in natural language, and GP-BO maintains a high-dimensional posterior. We provide the full procedure in Algorithm˜1.
@@
1 Initialize CMA-ES, $\mathcal{H}\leftarrow\emptyset$;
2 for $t=1,\dots,T$ do
3 if LLM turn (with probability $r$) then
4      Extract $\bm{\mu},\sigma,\mathbf{C}$ from CMA-ES;
5    $\mathbf{x}\leftarrow$ LLM($\bm{\mu},\sigma,\mathbf{C}$, $\mathcal{H}$, $\mathcal{S}$);
6
7 else
8    $\mathbf{x}\leftarrow$ CMA-ES.Propose();
9
10   end if
11 $y\leftarrow$ Evaluate($\mathbf{x}$);
12   CMA-ES.Update($\mathbf{x},y$);
13 $\mathcal{H}\leftarrow\mathcal{H}\cup\{(\mathbf{x},y)\}$;
14
15 end for
@@ 5.1 Classical methods outperform LLMs in fixed search spaces
Figure˜1 shows convergence curves against cumulative training wall-time for all 27B methods. Within the fixed search space, classical HPO methods consistently outperformed pure LLM-based approaches. The top methods by mean best val_bpb are Centaur (0.9763), TPE (0.9768), SMAC (0.9778), CMA-ES (0.9785), and Karpathy Agent (Code) (0.9814). The gap to the best fixed-space LLM method, LLAMBO (Paper) at 0.9862, is substantial, and several pure LLM methods performed worse than random search, indicating that restricting LLMs to a fixed HP search space does not leverage their strengths.
OOM avoidance matters more than search diversity. Table˜3 shows that the top methods all have OOM rates at or below 16%, while the bottom four exceed 36%. Karpathy Agent (14 HPs) has the lowest diversity by all metrics (Table˜3), converging to a narrow region early. LLAMBO (Paper) is one of the most diverse methods yet underperforms classical methods due to its 48% OOM rate. These OOM rates also illustrate the state-tracking limitation of small and mid-sized LLMs: LLAMBO (Paper) and LLAMBO (Optuna) observe full trial history yet produce OOM rates (48% and 61%) comparable to random search (56%), suggesting they fail to learn which regions of the search space cause memory failures. In contrast, CMA-ES and TPE maintain explicit optimization state and keep OOM rates at 16% and 11%. CMA-ES explores broadly but efficiently: it covers 802 grid cells with only 16% OOM, showing that covariance adaptation learns which directions are safe. However, CMA-ES also has the highest variance among top methods (std 0.0036 vs 0.0019 for TPE), with its best seed achieving the single best result in the benchmark (0.9741) while its worst seed is mediocre (0.9829).
@@ 5.2 Unconstrained code editing is viable but requires model scale
Karpathy Agent (Code), which directly edits training source code rather than operating in the fixed search space, is the only pure LLM method competitive with classical approaches. Given the simplicity of the setup and the use of a self-hosted open-weight model (Qwen3.5-27B), the gap to classical methods is smaller than one might expect, and stronger frontier models may close it further.
Scaling the LLM from 0.8B to 27B is essential for unconstrained code editing but provides no advantage for fixed-HP optimization. Figure˜2 shows that 0.8B is insufficient for unconstrained code editing (Karpathy Agent Code: 0.9910 vs 0.9814 with 27B), while fixed-HP methods saw no benefit from scaling (Karpathy Agent 14 HPs: 0.9904 vs 0.9908). All LLM methods used open-weight Qwen3.5, and the gap between unconstrained code editing and classical methods is already small; stronger frontier models may further improve LLM-based methods, particularly for unconstrained code editing where model capability directly determines the quality of code modifications. A preliminary comparison with Gemini 2.5 Flash (Comanici et al., 2025) (Section˜A.2) shows no improvement over Qwen3.5-27B, though stronger or larger frontier models may yield different results.
@@ 5.3 Hybrid optimization: best of both worlds
Centaur outperformed all methods including CMA-ES alone by using the LLM on only 30% of trials. As described in Section˜4, the LLM receives CMA-ES’s full internal state ($\bm{\mu}$, $\sigma$, $\mathbf{C}$), the top-5 configurations, and the last 20 trials, and almost always overrides CMA-ES’s proposal (100% for 27B, 95% for 0.8B). Despite constituting only 30% of trials, LLM trials contributed 25% of incumbent improvements (Section˜A.4), confirming that the LLM provides complementary value beyond what CMA-ES finds alone. Beyond improving the mean, Centaur substantially reduces CMA-ES’s cross-seed variance: std drops from 0.0036 for CMA-ES alone to 0.0005 for Centaur, suggesting that the LLM stabilizes the optimizer by injecting domain knowledge that prevents unfavorable seeds from drifting. Notably, Centaur [0.8B] outperformed Centaur [27B] (0.9766 vs 0.9763), suggesting that a cheap LLM suffices when paired with a strong classical optimizer. This stands in contrast to unconstrained code editing, where the 0.8B model is insufficient: when the classical optimizer handles the search trajectory, the LLM’s role reduces to refining promising candidates, a task that does not require the same level of model capability as generating code edits from scratch. We ablate the LLM ratio $r$ in Section˜A.6. For the 0.8B model, r=0.2r{=}0.2 achieves the best single-seed result (0.9735), slightly outperforming the default r=0.3r{=}0.3 (0.9776), while r=0.8r{=}0.8 degrades to 0.9789. For the 27B model, r=0.5r{=}0.5 performs best (0.9745), but r=0.8r{=}0.8 collapses to 0.9903, worse than CMA-ES alone. This confirms that CMA-ES should retain majority control of the optimization trajectory: the LLM is most effective as an occasional informed perturbation, not as the primary search driver.
@@ 6 Conclusion
We benchmarked classical, LLM-based, and hybrid HPO methods on small-scale language model training under the same budgets and constraints. Within the fixed, constrained search space, classical HPO methods such as CMA-ES and TPE consistently outperformed LLM-based agents, with OOM avoidance emerging as a stronger predictor of performance than search diversity. Restricting LLMs to a fixed hyperparameter search space does not leverage their strengths; the one setting where a pure LLM method competes with classical approaches is unconstrained code editing (Karpathy, 2025a), where the agent directly modifies training source code and narrows the gap to classical methods substantially despite using only a self-hosted open-weight model. However, this requires sufficient model scale: the 0.8B model is insufficient for unconstrained code editing, while it suffices for hybrid optimization. Centaur achieved the best result in our experiments by using the LLM on only 30% of trials, preserving fast convergence from CMA-ES while benefiting from LLM-informed suggestions, with the 0.8B variant outperforming the 27B variant.
Our study evaluated a single task with open-weight models (Qwen3.5 0.8B and 27B). Given that unconstrained code editing already nearly matches classical methods with these models, stronger frontier models may shift the balance further in favor of LLM-based approaches. While we automated hyperparameter selection, the search space ranges still require manual specification, introducing some human priors. Future work could benchmark unconstrained code editing with frontier models to determine whether increased model capability allows LLM agents to match or surpass hybrid and classical methods. Additionally, exploring other classical optimizers as the hybrid base and pairing CMA-ES with a code-editing LLM agent could allow the search space to co-evolve with the optimization trajectory.
Appendix A Additional Results

### 29 Mar 2026 v2

@@ Abstract
The autoresearch repository enables an LLM agent to search for optimal hyperparameter configurations on an unconstrained search space by editing the training code directly. Given a fixed compute budget and constraints, we use autoresearch as a testbed to compare classical hyperparameter optimization (HPO) algorithms against LLM-based methods on tuning the hyperparameters of a small language model. Within a fixed hyperparameter search space, classical HPO methods such as CMA-ES and TPE consistently outperform LLM-based agents. However, an LLM agent that directly edits training source code in an unconstrained search space narrows the gap to classical methods substantially despite using only a self-hosted open-weight 27B model. Methods that avoid out-of-memory failures outperform those with higher search diversity, suggesting that reliability matters more than exploration breadth. While small and mid-sized LLMs struggle to track optimization state across trials, classical methods lack domain knowledge. To bridge this gap, we introduce Centaur, a hybrid that shares CMA-ES’s internal state, including mean vector, step-size, and covariance matrix, with an LLM [^1] Named after the mythological half-human, half-horse hybrid: our method merges LLM reasoning with classical optimization.. Centaur achieves the best result in our experiments, with its 0.8B variant outperforming the 27B variant, suggesting that a cheap LLM suffices when paired with a strong classical optimizer. The 0.8B model is insufficient for unconstrained code editing but sufficient for hybrid optimization, while scaling to 27B provides no advantage for fixed search space methods. Preliminary experiments with the frontier model Gemini 3.1 Pro Preview do not close the gap to classical methods. Code is available at https://github.com/ferreirafabio/autoresearch-automl.
@@ 1 Introduction
To answer these questions, we benchmark 9 HPO methods, spanning 4 classical, 4 LLM-based, and 1 hybrid, on Karpathy’s autoresearch task. All methods operate under the same 24-hour GPU training budget with 3 seeds. To reduce human priors, we automatically extract 14 hyperparameters from the training script; while the ranges require some domain knowledge, the HP selection itself is automated, removing manual search space curation. Figure˜1 compares all 9 methods using 27B LLM variants against cumulative training wall-time, showing that classical methods find better configurations than LLM-based agents within the fixed search space. The exception is Karpathy Agent (Code), which directly edits training source code and is competitive with classical methods, even though the latter find a similar performing hyperparameter configuration ${\sim}4{\times}$ faster. All LLM methods use a self-hosted open-weight model (Qwen3.5-27B); preliminary experiments with the frontier model Gemini 3.1 Pro Preview (Section˜A.2) show competitive but not superior performance.
@@
  • We show that classical HPO outperforms LLM agents within a fixed search space. An LLM agent that directly edits code is more competitive but still falls short of the best classical methods.
@@
  • We analyze search diversity, OOM rates, and model scaling across all methods. Preliminary experiments with the frontier model Gemini 3.1 Pro Preview suggest that simply scaling the LLM optimizer does not close the gap to classical methods.
@@ 3 Experimental Setup
We ran each method for 24 hours with three seeds and report results against cumulative training wall-time. Methods with high OOM rates accumulated more trials due to fast failures (<30{<}30s per OOM); trial-number plots in the appendix are capped at 300 trials as no meaningful improvement occurs beyond that point. We report failed trials (OOM) as val_bpb=100.0{}=100.0, a finite penalty orders of magnitude worse than any valid result while remaining compatible with all surrogate models, so that optimizers learn to avoid infeasible regions.
Our main experiments use Qwen3.5 (Team, 2026) (0.8B and 27B) as the LLM optimizer, self-hosted via vLLM (Kwon et al., 2023) on the same GPU that trains the optimizee (the $\sim$}50M-parameter language model). We additionally run frontier model experiments with Gemini 2.5 Flash (Comanici et al., 2025), 3.1 Flash-Lite, and 3.1 Pro Preview (Google DeepMind, 2026) via the Gemini API. To ensure equal resource allocation, we capped training VRAM to 80 GB for all methods, comparable to the H100 used by Karpathy (Karpathy, 2025a), and reserved the remaining memory for the vLLM server when using self-hosted models. We disabled Qwen3.5’s thinking mode and sampled with temperature 0.7, limiting outputs to 2048 tokens for fixed-HP methods and 16384 tokens for code editing. All prompt templates are available in our repository.
@@ 4 Centaur: CMA-ES Guided LLM Optimization
Centaur shares CMA-ES’s full internal state with the LLM on a fraction of trials. On every trial, CMA-ES proposes a candidate configuration from its multivariate Gaussian, parameterized by mean vector $\bm{\mu}$, step-size $\sigma$, and covariance matrix $\mathbf{C}$. On 30% of trials, the LLM receives CMA-ES’s proposal along with $\bm{\mu}$, $\sigma$, $\mathbf{C}$, the top-5 configurations, and the last 20 trials.
@@
Initialize CMA-ES, history $\mathcal{H}\leftarrow\emptyset$
for $t=1,\dots,T$ do
  if LLM turn (with probability $r$) then
     Extract $\bm{\mu},\sigma,\mathbf{C}$ from CMA-ES
     $\mathbf{x}\leftarrow$ LLM($\bm{\mu},\sigma,\mathbf{C}$, $\mathcal{H}$, $\mathcal{S}$)
 else
     $\mathbf{x}\leftarrow$ CMA-ES.Propose()
 $y\leftarrow$ Evaluate($\mathbf{x}$)
  CMA-ES.Update($\mathbf{x},y$)
  $\mathcal{H}\leftarrow\mathcal{H}\cup\{(\mathbf{x},y)\}$
@@
The LLM may override the proposal, and in practice does so in nearly all cases: 100% of the time with the 27B model and 95% with the 0.8B model. Crucially, CMA-ES updates its internal state from all trial results, including those where the LLM overrode its proposal, so the optimizer continuously learns from the full trajectory.
We chose CMA-ES because its internal state is particularly interpretable for LLM communication: the mean is a concrete configuration, $\sigma$ is a single scalar, and $\mathbf{C}$ is a labeled matrix. In contrast, TPE maintains density estimators that are difficult to summarize in natural language, and GP-BO maintains a high-dimensional posterior.
@@ 5.1 Classical methods outperform LLMs in fixed search spaces
Figure˜1 shows convergence against cumulative training wall-time for all 27B methods. Within the fixed search space, classical HPO methods consistently outperformed pure LLM-based approaches. The top methods by mean best val_bpb are Centaur (0.9763), TPE (0.9768), SMAC (0.9778), CMA-ES (0.9785), and Karpathy Agent (Code) (0.9814). The gap to the best fixed-space LLM method (LLAMBO (Paper)) is significant, and several pure LLM methods performed worse than random search, indicating that restricting LLMs to a fixed HP search space does not leverage their strengths.
Moreover, OOM avoidance matters more than search diversity: Table˜3 shows that the top methods all have OOM rates at or below 16%, while the bottom four exceed 36%. Karpathy Agent (14 HPs) has the lowest diversity by all metrics, converging to a narrow region early. LLAMBO (Paper) is one of the most diverse methods yet underperforms classical methods due to its 48% OOM rate. These OOM rates also illustrate the state-tracking limitation of small and mid-sized LLMs: LLAMBO (Paper) and LLAMBO (Optuna) observe full trial history yet produce OOM rates (48% and 61%) comparable to random search (56%), suggesting they fail to learn which regions of the search space cause memory failures. In contrast, CMA-ES and TPE maintain explicit optimization state and keep OOM rates at 16% and 11%, showing that covariance adaptation and density estimation learn which regions are safe. However, CMA-ES also has the highest variance among top methods (std 0.0036 vs 0.0019 for TPE), with its best seed achieving the single best result in the benchmark (0.9741) while its worst seed is mediocre (0.9829).
@@ 5.2 Unconstrained code editing is viable but requires model scale
Karpathy Agent (Code), which directly edits training source code rather than operating in the fixed search space, is the only pure LLM method competitive with classical approaches. Given the simplicity of the setup and the use of a self-hosted open-weight model (Qwen3.5-27B), the gap to classical methods is smaller than one might expect.
Scaling the LLM from 0.8B to 27B is essential for unconstrained code editing but provides no advantage for fixed-HP optimization, as can be seen in Figure˜2: 0.8B is insufficient for unconstrained code editing (Karpathy Agent Code: 0.9910 vs 0.9814 with 27B), while fixed-HP methods saw no benefit from scaling (Karpathy Agent 14 HPs: 0.9904 vs 0.9908). All LLM methods in Figures˜1 and 2 used open-weight Qwen3.5. Preliminary comparisons with Gemini 2.5 Flash (Comanici et al., 2025) and Gemini 3.1 Flash-Lite (Google DeepMind, 2026) as the LLM optimizer (Section˜A.2) show no improvement over Qwen3.5-27B. Gemini 3.1 Flash-Lite fails to generate valid code in 87–94% of trials for unconstrained editing, but performs comparably to Qwen3.5-0.8B in the fixed search space, confirming that code editing is where model capability matters most. Gemini 3.1 Pro Preview (Figure˜3) is competitive with Qwen3.5-27B for both unconstrained code editing and Centaur, with lower OOM rates for code editing (5% vs 12%), but does not close the gap to the best classical methods. Gemini 3.1 Pro Preview results are based on a single seed; with the frontier models tested so far, we do not see evidence that simply scaling the LLM optimizer shifts the balance in favor of LLM-based approaches.
Figure 3: Gemini 3.1 Pro Preview vs Qwen3.5-27B for Karpathy Agent (Code) and Centaur. Solid: Gemini (single seed, due to API cost constraints), dashed: Qwen (3 seeds). TPE shown as reference.
@@ 5.3 Hybrid optimization: best of both worlds
Centaur outperformed all methods including CMA-ES alone by using the LLM on only 30% of trials. As described in Section˜4, the LLM receives CMA-ES’s full internal state ($\bm{\mu}$, $\sigma$, $\mathbf{C}$), the top-5 configurations, and the last 20 trials, and almost always overrides CMA-ES’s proposal (100% for 27B, 95% for 0.8B). Despite constituting only 30% of trials, LLM trials contributed 25% of incumbent improvements (Section˜A.5), confirming that the LLM provides complementary value beyond what CMA-ES finds alone. Beyond improving the mean, Centaur substantially reduces CMA-ES’s cross-seed variance: std drops from 0.0036 for CMA-ES alone to 0.0005 for Centaur, suggesting that the LLM stabilizes the optimizer by injecting domain knowledge that prevents unfavorable seeds from drifting. Notably, Centaur [0.8B] outperformed Centaur [27B], suggesting that a cheap LLM suffices when paired with a strong classical optimizer. This stands in contrast to unconstrained code editing, where the 0.8B model is insufficient: when the classical optimizer handles the search trajectory, the LLM’s role appears to reduce to refining promising candidates, a task that does not require the same level of model capability as generating code edits from scratch. We ablate the LLM ratio $r$ (see Appendix Tables˜4 and 8). For the 0.8B model, r=0.2r{=}0.2 achieves the best result (0.9735), slightly outperforming the default r=0.3r{=}0.3 (0.9776), while r=0.8r{=}0.8 degrades to 0.9789. For the 27B model, r=0.5r{=}0.5 performs best (0.9745), but r=0.8r{=}0.8 collapses to 0.9903, which is worse than CMA-ES alone. This confirms that CMA-ES should retain majority control of the optimization trajectory: the LLM is most effective as an occasional informed perturbation, not as the primary search driver.
@@ 6 Conclusion
We benchmarked classical, LLM-based, and hybrid HPO methods on small-scale language model training under the same budgets and constraints. Within the fixed, constrained search space, classical HPO methods such as CMA-ES and TPE consistently outperformed LLM-based agents, with OOM avoidance emerging as a stronger predictor of performance than search diversity. Restricting LLMs to a fixed hyperparameter search space does not leverage their strengths; the one setting where a pure LLM method competes with classical approaches is unconstrained code editing (Karpathy, 2025a), where the agent directly modifies training source code and narrows the gap to classical methods substantially despite using only a self-hosted open-weight model. However, this requires sufficient model scale: the 0.8B model is insufficient for unconstrained code editing, while it suffices for hybrid optimization. Centaur achieved the best result in our experiments by using the LLM on only 30% of trials, preserving fast convergence from CMA-ES while benefiting from LLM-informed suggestions.
Our study evaluated a single task with open-weight models (Qwen3.5 0.8B and 27B). Preliminary single-seed experiments with frontier model Gemini 3.1 Pro Preview shows competitive but not clearly superior performance to Qwen3.5-27B (Section˜A.2). More experiments and other models are needed to determine whether frontier models can close the remaining gap to classical methods. Also, all LLM methods in this benchmark operate without tools or external knowledge retrieval. Equipping LLM optimizers with agentic capabilities such as paper search, documentation lookup, or code analysis tools is a promising direction. Additionally, exploring other classical optimizers as the hybrid base and pairing CMA-ES with a code-editing LLM agent could allow the search space to co-evolve with the optimization trajectory.
Acknowledgments and Disclosure of Funding
This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 539134284, through EFRE (FEIH_2698644) and the state of Baden-Württemberg. We acknowledge funding by the European Union (via ERC Consolidator Grant DeepLearning 2.0, grant no. 101045765). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Frank Hutter acknowledges the financial support of the Hector Foundation.
[Uncaptioned image][Uncaptioned image]
Appendix A Appendix